• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 16:09:17 on October24,2005

    Entry number 157890

    keyword=day shift summary

    Day shift summary:

    shift workers: J. P. Chen (SL), X. Jiang (TO), S. Fuchs (3rd).

    8:00 production run continue
    8:50 Compton has been running with ROC5 down. Xiaodong stopped run and tried reboot, failed. Hachem is working on it.
    9:02 asked for IHWP IN, did flip.

    9:45, beam off, pass change for Hall B
    9:45 Compton is back.

    11:00 controlled access for cryo work on the right dipole.
    11:30 right magnets went off, cryo level low and then went to zero.
    11:40 right septum temperature went to high (>8K). informed Ed.
    11:40 Per Jack's request, pushed reset (green) button for the right
    spectrometer control twice.
    12:00 left cryo level low, Q1 and Dipole went to zero.
    13:30 MCC started to record Hall downtime for magnet repair.
    14:30 per Jack's request, ramped left Q2/Q3 to zero.
    15:50 hall controlled access ended. request beam permit.
    16:00 beam back, 5 uA.
    16:00 spectrometer magnets all set except right dipole, which is still ramping up.


    run 5856, 7:33-8:29, 100K events, 38K OK pairs, IHWP OUT.
    run 5857, 8:29-9:02, 57K, 20K OK pairs, IHWP OUT

    run 5858, 9:04-9:45, 73K, 27K OK pairs, IHWP IN.

    run 5859-5861 cosmic runs.